[figure: 9.7 cm]

TEMPERATURE DEPENDENCE OF MAGNETISATION REVERSAL AND GMR IN SPIN VALVE STRUCTURES

F. Stobiecki^{a}, T. Stobiecki^{b}, B. Ocker^{c}, W. Maass^{c}, W. Powroźnik^{b}, A. Paja^{b}, C. Loch^{d}, and K. Röll^{d} aInstitute of Molecular Physics, Polish Academy of Sciences, Poznań, Poland

bDepartment of Electronics, University of Mining and Metallurgy, Kraków, Poland

cBalzers Process Systems GmbH, Alzenau, Germany dDepartment of Physics, University of Kassel, Kassel, Germany

Spin valve structures consisting of Si/Al_{2}O_{3}/Ta/Co-4.4 nm/Cu-2.2 nm/Co-4.4 nm/ FeMn/Ta/Cu deposited by sputtering in different conditions show essential differences in GMR effect (GMR was about 6% for the worst and 11% for the best samples at RT). Two contributions should be taken into consideration in explaining this effect: i) differences in ferromagnetic coupling between Co layers (for a strong coupling antiparallel alignment of magnetisation of Co layers is not achieved) [1], ii) differences in resisitivity and mean free path of electrons [2]. The temperature measurements of R(H) and m(H) dependences (R - resistance, m - magnetic moment, H - magnetic field) enable us to separate both contributions to GMR effect and to explain which kind of ferromagnetic coupling (RKKY - like magnetostatic coupling or pinholes) is dominating in particular spin valves.

[1] Th.G.S. Rijks et al., J. Appl. Phys. 76 (1994) 1092

[2] B. Dieny, J. Magn. Magn. Mater. 136 (1994) 335

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : F. Stobiecki

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences PL 60-179 Poznań ul. Smoluchowskiego 17, Poland

Email address : stfeliks@ifmpan.poznan.pl